OpenStudy (anonymous):

Rationalize the denominator of \[\frac{sqrt{x^2}}{sqrt{x}+1}\]

OpenStudy (anonymous):

\[\frac{\sqrt{x^2}}{\sqrt{x}+1}\]

OpenStudy (anonymous):

I got \[\frac{x^3-x}{x-1}\]

OpenStudy (anonymous):

\[\frac{\sqrt{x^2}}{\sqrt{x}+1} = \frac{x}{\sqrt{x}+1}=\frac{x(\sqrt{x}-1)}{(\sqrt{x}+1)(\sqrt{x}-1)}=\frac{x\sqrt{x}-x}{x-1}\]

OpenStudy (anonymous):

I see what I did wrong thanks @RolyPoly

OpenStudy (anonymous):

Note that \[\sqrt{x^2} = x\]

OpenStudy (anonymous):

Okay, I noted that :). Thank You once again!

OpenStudy (anonymous):

welcome